A boat can travel 9 miles upstream in the same amount of time it takes to travel 11 miles downstream. if the current of the river is 3 miles per hour, complete the chart below and use it to find the speed of the boat in still water
distance = rate * Time
upstream 9 r-3
downstream 11 r+3

Found 2 solutions by solver91311, nyc_function:
Answer by solver91311(17077)



The time is the same for both the upstream and downstream trips ("in the same amount of time"), so just call it .

Since


we can also say



So for the upstream trip:



And for the downstream trip:



But , so we can say:



Just cross-multiply and solve for

John


Answer by nyc_function(2733)

We need to find the time for the up and down trip.
We use the following equation twice to find the time going up and down: d/r = t

Going Up:
9 = (r + 3)t
9/(r + 3) = t
Going Down:
11 = (r - 3)t
11/(r - 3) = t

I will set up the equation and you can finish.

(r + 3)(9/(r + 3) = (r - 3)(11/(r - 3)
Can you finish now?